### v2

@@ 4.3 Awareness and Education
Awareness of reproducibility issues and available training/education to support reproducibility can be a powerful driver for ML reproducibility [77]).
@@
To enhance awareness and establish a minimum of reproducibility standards, the policies of scientific journals are considered an influence factor. A number of journals already mandate data and/or code availability for publication [59, 58, 31]. However, to address issues such as result manipulation, more extensive journal participation is needed to for instance introduce preregistration, where researchers register their research intentions for future publication. This approach ensures credibility by separating the research plan from experimental outcomes [72, 54], thereby reducing spin practices, HARKing, and p-hacking [30]. The ACM TORS (Transactions on Recommender Systems) journal exemplifies this by allowing preregistration and publishing “reproducibility papers” dedicated to reproduction studies and enhancing reproducibility tools. Apart from that, various initiatives have been launched to raise awareness of reproducibility issues. A few examples are the following:
  • The ReScience journal publishes peer-reviewed papers discussing attempts to reproduce original publications. These reproductions are published on GitHub [^19] https://github.com/, being available to other researchers [66].
@@
  • Reproducibility challenges, where several researchers try to reproduce many recent publications in parallel, are being held frequently. These challenges allow for an analysis of the success rate of reproduction and can be used to evaluate progress over multiple years [59]. Additionally, conferences such as the European Conference on Information Retrieval (ECIR), provides special reproducibility tracks, in which researchers are encouraged to reproduce existing papers and build upon their results (e.g., [43, 50]).
  • The ACM has put together a new emerging interest group on reproducibility [^21] https://reproducibility.acm.org/. The main goals are to (i) contribute to the development of reproducibility standards, practices and policies, (ii) promote the development and evaluation of tools, techniques, and methodologies, and (iii) encourage best practices.
@@ 5 Mapping Drivers to Barriers
Our Drivers-Barriers-Matrix shows that there are often multiple drivers for the same barrier. Consequently, there are also several possible solutions for a barrier or different aspects of a barrier. The mapping allows to quickly assess which drivers address the different barriers and which barriers have a higher or lower number of drivers associated with them. It underlines the need for context-dependent approaches instead of "one-size-fits-all" solutions, as the proper selection of a suitable driver depends on the specific conditions and existing barriers relevant to any ML application. Intersections between drivers and barriers will be described in more detail below.
@@
The major drivers for completeness and report quality are guidelines and checklists. Guidelines provide best practices to adopt in order to achieve reproducible ML research. Furthermore, many checklists exist that comprehensively state the different pitfalls and provide information on how they can be avoided. Researchers can use them to ensure their research meets the desired standards. Furthermore, some of these checklists and guidelines are enforced by journals, such that research will only be published if certain criteria are met. In comparison, spin practices are not as easily identifiable. In this case, the discussion within the research community centers around removing the incentives for inflating research results. A particularly effective driver for this is preregistration as an example for publication policies and initiatives, where researchers submit research proposals before conducting an experiment. If accepted, the research will be published, regardless of the outcome, thereby minimizing spin practices.
Figure 2: Drivers-Barriers-Matrix. We map the 9 drivers to the 9 barriers identified in this paper. The colored boxes show that a specific driver is applicable to a specific barrier.
@@
Code sharing is essential to reproducibility, which makes limited access to code a significant barrier in the field. However, it is often neglected as the process is not trivial. To make shared code useful to the scientific community, it is necessary to share, in addition to the source code, the information about the entire software setup and dependencies, including software versions and hardware configurations. To assist with this, researchers may consider running their code in hosting services or virtualization environments, which we identified as drivers for code sharing. Both have similar advantages, i.e., they can easily be shared and made public for other researchers to use. As a consequence, it will give reproducers immediate access to code, including the complete configuration setup, such as dependencies and versions. Hosting services are the quicker and easier way of achieving this; however, they may be subject to different resource limits. Virtualization (e.g., VMs or containers) is more difficult to set up but offers more flexibility and is not externally (e.g., by a provider) restricted in capabilities and resources. Furthermore, tools and platforms can be drivers for reproducibility. A lot of ML tools provide capabilities for code versioning or other features, which are key to reproducibility. One example is dToolAI [32], which automatically logs the supplemental information of the code, i.e., metadata, hyperparameters, and more, which are essential for ML reproducibility.
@@
Data-related barriers are a severe obstacle to ML reproducibility due to the research fields’ data-driven nature, where limited access to data forms a major challenge. Privacy concerns are among the crucial arguments that cause hesitation in sharing data. The need for data privacy is evident, especially in the Biomedical fields, which deal with patients’ electronic health records. Nevertheless, it increases reproducibility issues in ML-based science and, thus, delays technological progress within these domains. However, there are several approaches that aim to meet the requirements of sharing sensitive data: Privacy-preserving technologies allow reproducers to train ML models on private data without actually possessing the data. This way, reproduction becomes possible without violating potential privacy regulations. Other than that, the use of standardized datasets and evaluation can support issues in regard to dataset meta-information, including the specification of train-test splits and data provenance. Once again, tools and platforms can assist with data versioning, and numerous guidelines and checklists have been proposed to address the provenance of data. These guidelines and checklists are designed to help researchers to avoid common pitfalls. Current initiatives are supported by journals that more frequently require data to be shared as part of a publication.
Concerning methodological errors associated with the data, data leakage is a major issue, which can, for instance, be mitigated using standardized datasets and evaluation. Other drivers to solve data leakage are model info sheets and model cards, which are provided as supplemental information to a published dataset. Even though there are some limitations to model info sheets, they are capable of detecting all types of data leakage.
@@
Bias is another methodological error, leading to irreproducible results. This is because the biased data usually does not generalize well to problems outside the experimental setup of a specific ML study. Bias has been an important source of concern, e.g., in Biomedical fields. Effects thereof can again be minimized using standardized datasets and evaluation or specific guidelines and checklists, e.g., ROBUST-ML [2].
@@
If an ML experiment is shared entirely and code and data are available, i.e., reproducibility type R4 Experiment, there are still three barriers, which can lead to irreproducible results. Inherent non-determinism arises from the different sources of randomness in ML, and makes it difficult to achieve repeatable results, even on the same machine. There are, however, methods to manage the sources of randomness, such as fixed random seeds and deterministic implementations, while comprehensively mitigating all sources of randomness is still a very challenging endeavor.
Another barrier is described as environmental differences, which has two main issues associated with it, i.e., software differences and hardware differences. Both types of differences can be avoided by using either hosting services or virtualization; constraints can be assumed to be similar to the barrier of limited access to code. Limited access to computational resources constitutes another barrier to ML reproducibility identified in this work. The issue is particularly problematic for research using LLMs because of their need for extensive computational resources in training and reproduction. Hosting services offer a solution, providing access to pre-trained models and allowing researchers to directly access and run respective models on-site.
@@ 6 Conclusion
In this paper, we examined the barriers and drivers associated with the four types of ML reproducibility as outlined by Gundersen et al. (description, data, code, and experiment) [26], specifically in the contexts of Computer Science and Biomedical research. We then synthesized this into a Drivers-Barriers-Matrix to summarize and illustrate which drivers are feasible solutions for the various barriers. We observe that the barriers to ML reproducibility can be addressed through three primary drivers: technology-driven solutions, procedural improvements, and enhanced awareness and education. It is important to highlight that, in theory, awareness and education can complement the other drivers and serve as a foundational step toward overcoming any reproducibility-related challenge. Further, the rise of AutoML tools for ML development and ML tasks performed by domain experts, (potentially) not having in-depth computer or data science knowledge, could become another barrier for reproducibility. However, we are confident that our findings will also help in this regard and that the topic likely requires collaborative efforts over the coming years.
We hope that our paper provides practical guidance and orientation for researchers employing ML, and clarifies the current state of play. Of course, in such a dynamic and fast-paced research area, this only opens a series of further questions and avenues for exploration. We recommend further investigation of the various issues and potential solutions laid out here. We would also encourage further investigation into the potential role of platforms [30] or foundation models [35] in further exacerbating or alleviating these issues.
6.0.1 Acknowledgements.
This research is supported by the Horizon Europe project TIER2 (GA: 101094817), and the FFG COMET program.

### v3

@@ 4.3 Awareness and Education
Awareness of reproducibility issues and available training/education to support reproducibility can be a powerful driver for ML reproducibility [117].
@@
To enhance awareness and establish a minimum of reproducibility standards, the policies of scientific journals are considered an influencing factor. A number of journals already mandate data and/or code availability for publication [93, 92, 50]. However, to address issues such as result manipulation, more extensive journal participation is needed to, for instance, introduce preregistration where researchers register their research intentions for future publication. This approach ensures credibility by separating the research plan from experimental outcomes [112, 87], thereby reducing spin practices, HARKing, and p-hacking [46]. The ACM TORS (Transactions on Recommender Systems) journal exemplifies this by allowing preregistration and publishing “reproducibility papers” dedicated to reproduction studies and enhancing reproducibility tools. Apart from that, various initiatives have been launched to raise awareness of reproducibility issues. A few examples are the following:
  • The ReScience journal publishes peer-reviewed papers discussing attempts to reproduce original publications. These reproductions are published on GitHub [^24] https://github.com/ and available to other researchers [103].
@@
  • Reproducibility challenges, where several researchers try to reproduce many recent publications in parallel, are being held frequently. These challenges allow for an analysis of the success rate of reproduction and can be used to evaluate progress over multiple years [93]. Additionally, conferences such as the European Conference on Information Retrieval (ECIR), provides special reproducibility tracks, in which researchers are encouraged to reproduce existing papers and build upon their results (e.g., [69, 83, 68]).
  • The ACM has convened a new emerging interest group on reproducibility [^26] https://reproducibility.acm.org/. The main goals are to (i) contribute to the development of reproducibility standards, practices and policies, (ii) promote the development and evaluation of tools and methodologies, and (iii) encourage best practices.
@@
As also indicated by research in information retrieval and recommender systems, increased awareness and education in the form of publication policies and initiatives can address reproducibility issues by emphasizing robust experimental practices, methodological rigor, and the development of shared resources among the different actors identified, i.e., students, educators, scholars, practitioners and decision-makers [12].
@@ 5 Mapping Drivers to Barriers
Figure 2: Drivers-Barriers-Matrix. We map the 9 drivers to the 9 barriers identified in this paper. The colored boxes show that a specific driver is applicable to a specific barrier. We argue that drivers related to awareness and education are, in general, applicable to address all barriers.
@@
Our Drivers-Barriers-Matrix shows that there are often multiple drivers for the same barrier. Consequently, there are also several possible solutions for a barrier or different aspects of a barrier. The mapping allows us to quickly assess which drivers address the different barriers and which barriers have a higher or lower number of drivers associated with them. It underlines the need for context-dependent approaches instead of “one-size-fits-all” solutions, as the proper selection of a suitable driver depends on the specific conditions and existing barriers relevant to any ML application. We describe intersections between drivers and barriers in more detail in the following and close this section with an overview of strengths and potential limitations of the identified drivers.
@@
The major drivers for completeness and report quality are guidelines and checklists. Guidelines provide best practices to adopt in order to achieve reproducible ML research. Furthermore, many checklists exist that comprehensively state the different pitfalls and provide information on how they can be avoided. Researchers can use them to ensure their research meets the desired standards. Furthermore, some of these checklists and guidelines are enforced by journals, such that research will only be published if certain criteria are met. In comparison, spin practices are not as easily identifiable. In this case, the discussion within the research community centers around removing the incentives for inflating research results. A particularly effective driver for this is preregistration as an example for publication policies and initiatives, where researchers submit research objectives and methods for review before conducting the research. If accepted, the research will be published regardless of the outcome (i.e., whether results are positive, negative or null), thereby minimizing spin practices.
@@
Code sharing is essential to reproducibility, which makes limited access to code a significant barrier in the field. However, it is often neglected as the process is not trivial. To make shared code useful to the scientific community, it is necessary to share, in addition to the source code, the information about the entire software setup and dependencies, including software versions and hardware configurations. To assist with this, researchers may consider running their code in hosting services or virtualization environments, which we identified as drivers for code sharing. Both have similar advantages, i.e., they can easily be shared and made public for other researchers to use. As a consequence, it will give reproducers immediate access to code, including the complete configuration setup, such as dependencies and versions. Hosting services are a quicker and easier way of achieving this; however, they may be subject to different resource limits. Virtualization (e.g., VMs or containers) is more difficult to set up but offers more flexibility and is not externally (e.g., by a provider) restricted in capabilities and resources. Furthermore, tools and platforms can be drivers for reproducibility. A lot of ML tools provide capabilities for code versioning or other features, which are key to reproducibility. One example is dToolAI [52], which automatically logs the supplemental information of the code, i.e., metadata, hyperparameters, and more, which are essential for ML reproducibility.
@@
Data-related barriers are a severe obstacle to ML reproducibility due to the research fields’ data-driven nature, where limited access to data forms a major challenge. Privacy concerns are among the crucial arguments that cause hesitation in sharing data. The need for data privacy is evident, especially in biomedical fields, which deal with patients’ electronic health records. Nevertheless, it increases reproducibility issues in ML-based science and, thus, delays technological progress within these domains. However, there are several approaches that aim to meet the requirements of sharing sensitive data: Privacy-preserving technologies allow reproducers to train ML models on private data without actually possessing the data. This way, reproduction becomes possible without violating potential privacy regulations. Other than that, the use of standardized datasets and evaluation can support issues in regard to dataset meta-information, including the specification of train-test splits and data provenance. Once again, tools and platforms can assist with data versioning, and numerous guidelines and checklists have been proposed to address the provenance of data. These guidelines and checklists are designed to help researchers to avoid common pitfalls. Current initiatives are supported by journals that more frequently require data to be shared as part of a publication.
@@
Concerning methodological errors associated with the data, data leakage is a major issue, which can, for instance, be mitigated using standardized datasets and evaluation. Other drivers to solve data leakage are model info sheets and model cards, which are provided as supplemental information to a published dataset. Even though there are some limitations to model info sheets, they are capable of detecting all types of data leakage. Bias is another methodological error, leading to irreproducible results. This is because the biased data usually does not generalize well to problems outside the experimental setup of a specific ML study. Bias has been an important source of concern, e.g., in biomedical fields. Effects thereof can again be minimized using standardized datasets and evaluation or specific guidelines and checklists, e.g., ROBUST-ML [3].
Table 1: Comparison of strengths and weaknesses of our identified drivers.
Driver Strengths Potential Limitations
Hosting Services Facilitates sharing of models, code, and datasets; increases accessibility. Out-of-the-box reproducibility is not yet provided and there are limits to the available compute.
Virtualization Enables environment replication (e.g., Docker, virtual machines); resolves dependency issues. Requires technical expertise; may introduce overhead for simple experiments.
Managing sources of randomness Critical for deterministic outcomes; has the potential to reduce and even eliminate variances across multiple runs. Can be hard to implement consistently across frameworks; only leads to point estimates of performance.
Privacy-preserving technologies Expands access to sensitive datasets without compromising privacy. Still an emerging field; performance trade-offs can make widespread adoption slower.
Tools and platforms Can streamline reproducibility practices and automatically acquire reproducibility artifacts. Out-of-the-box reproducibility is not yet provided, and fragmentation of tools can lead to siloed solutions rather than unified workflows.
Standardized datasets and evaluation Provides consistency and comparability for results across studies. May not generalize well to niche or domain-specific problems and can be subject to privacy concerns.
Guidelines, checklists Promotes best practices through structured processes (e.g., reproducibility checklists). Compliance can be time-consuming and may not be enforced consistently.
Model info sheets and model cards Improves transparency around model design and intended use. Adoption is still limited; requires effort to standardize and maintain across the community.
Publication policies, initiatives Drive cultural change by incentivizing openness (e.g., benchmarks, competitions). Impact depends on community participation and is a slow process in general.
@@
If an ML experiment is shared entirely and code and data are available, i.e., reproducibility type R4 Experiment, there are still three barriers, which can lead to irreproducible results. Inherent nondeterminism arises from the different sources of randomness in ML, and makes it difficult to achieve repeatable results, even on the same machine. There are, however, methods to manage the sources of randomness, such as fixed random seeds and deterministic implementations, while comprehensively mitigating all sources of randomness is still a very challenging endeavor.
Another barrier is described as environmental differences, which has two main issues associated with it, i.e., software differences and hardware differences. Both types of differences can be avoided by using either hosting services or virtualization; constraints can be assumed to be similar to the barrier of limited access to code. Limited access to computational resources constitutes another barrier to ML reproducibility identified in this work. The issue is particularly problematic for research using LLMs because of their need for extensive computational resources in training and reproduction. Hosting services offer a solution, providing access to pre-trained models and allowing researchers to directly access and run respective models on-site. Finally, Table 1 gives an overview of strengths and potential limitations of the identified drivers. As we can see, the choice of using a particular driver strongly depends on the given use and to what extent potential limitations are applicable for the use case.
@@ 6 Conclusion
In this paper, we examined the barriers and drivers associated with the four types of ML reproducibility as outlined by Gundersen et al. (description, data, code, and experiment) [43], specifically in the cases of computer science and biomedical research. We synthesized our findings into a Drivers-Barriers-Matrix to summarize and illustrate which drivers are feasible solutions to the various barriers. We observe that the barriers to ML reproducibility can be addressed through three kinds of drivers: technology-driven solutions, procedural improvements, and enhanced awareness and education. It is important to highlight that, in theory, awareness and education can complement the other drivers and serve as a foundational basis for overcoming reproducibility-related challenges.
One of the main issues hindering reproducibility in research appears to be rooted in the cultural aspects of research communities. As argued by [22, 12], the current incentives for conducting reproducible research are limited, and open research is often regarded as an unrewarded additional effort. Consequently, there is a lack of training and insufficient funding to cover the additional time and resources required by researchers. Notably, the lack of funding also impacts the ability to perform quality checks during and after the publication process. Therefore, we strongly believe that the way forward towards ML reproducibility is rooted in better education and more awareness of this topic among all involved stakeholders, e.g., students, educators, researchers, publishers, and policymakers. This, combined with the other tools and drivers described in this paper, could lead to more reproducible ML pipelines and, with this, more robust findings.
From a more technical perspective, the rise of AutoML tools for ML development and ML tasks performed by domain experts, (potentially) not having in-depth computer or data science knowledge, could pose another barrier to reproducibility [48]. We thus believe future work should address the increasing use of AutoML tools for AI development in research ([48]) among non-computer or data science experts. While these easy-to-use tools can standardize ML workflows by default and include documentation features, domain experts often lack the necessary expertise to recognize potential problems associated with ML, such as biased or imbalanced data. Thus, research on reproducibility should emphasize this challenge and aim to establish standards and guidelines for the use of No and Low Code ML tools in research, as well as the training required for their responsible application.
In summary, we hope that our paper provides practical guidance and orientation for researchers employing ML and clarifies the current state of play. Of course, in such a dynamic and fast-paced research area, this discussion opens up a series of further questions and avenues for exploration. We recommend further investigation of the various issues and potential solutions laid out here. We would also encourage further investigation into the potential role of platforms [46] or foundation models [55] in further exacerbating or alleviating these challenges.
Acknowledgements. This research is supported by the Horizon Europe project TIER2 (GA: 101094817), and the FFG COMET program.